OpenStudy (anonymous):

Kerri correctly calculated the distance between point S (–3428) and point T (286) on a number line. What value did Kerri find for the distance? A. –3714 B. –3142 C. 3142 D. 3714

OpenStudy (readergirl12):

To get to 0 from -3428 all you have to do is add the same number to itself. So like if I had -2 and I added 2 to it I'd get 0. So If I say -3428 + 3428 = 0 and 0 + 286 = 286 then 3428 + 286 = the distance.

OpenStudy (anonymous):

so the answer is a or d

OpenStudy (readergirl12):

Its B :P

OpenStudy (anonymous):

oh I was off but are you sure

OpenStudy (readergirl12):

I used my calculator x3

OpenStudy (readergirl12):

Im sure its B

OpenStudy (readergirl12):

Well No I just redid it Its D

OpenStudy (anonymous):

lmao really lol ok thanks

OpenStudy (readergirl12):

No probs ^-^